### v1

# Semiregular abstract polyhedra with trivial facet stabilizer

Abstract
Abstract polytopes generalize the face lattice of convex polytopes. An (abstract) polytope is semiregular if its facets are regular and its automorphism group acts transitively on its vertices. In this paper we construct semiregular, facet-transitive polyhedra with trivial facet stabilizer, showing that semiregular abstract polyhedra can have an unbounded number of flag orbits, while having as little as one facet orbit. We interpret this construction in terms of operations applied to high rank regular and chiral polytopes, and we see how this same operations help us construct alternating semiregular polyhedra. Finally, we give an idea to generalize this construction giving examples in higher ranks.
1 Introduction
Abstract polytopes are combinatorial objects that generalize the face lattice of convex polytopes and other geometric objects. Just as with geometric polytopes, the main approach to study abstract polytopes has been through their symmetry. A flag is typically defined as a set having one face of each rank, and satisfying that all those faces are incident to each other. One can prove that the automorphism group of an abstract polytope acts semiregularly on its flags, and therefore, one can measure how symmetric a polytope is by counting its flag orbits: the less flag orbits the more symmetrical. Regular polytopes are those with exactly one flag orbit, and are by far the most studied class of abstract polytope, having a whole book [10] dedicated to them.
When studying other classes of polytopes, some have studied polytopes with a small number of flag orbits (see for example [6], [12], [16], [17]), but some classes where the number of flag orbits is not restricted have been studied too. Such is the case of semiregular polytopes: those with regular facets and where the automorphism group acts transitively on the vertices. Despite the number of flag orbits being unbounded, most of the research on semiregular polytopes has been focused on constructions where the result has only one or two flag orbits (see [14], for example). This is due to the fact that the examples constructed are hereditary: every symmetry of a facet induces a symmetry of the whole polytope.
In this paper we construct semiregular polyhedra that are face-transitive, that are the opposite of hereditary: No symmetry of a facet induces a symmetry of the whole polyhedron, in other words, our examples have trivial facet stabilizer. This implies that despite having one orbit on vertices and one orbit on faces, the number of orbits on flags in unbounded in our construction.
For our construction is convinient to think of polytopes as a special class of maniplexes: graphs with a proper coloring of their edges satisfying certain conditions. Thinking of polytopes in terms of graphs allows us to use voltage graphs to construct them, and thanks to the results of [7] and [1]. we know conditions on the voltage assignments that tell us when the graphs we obtain represent a polytope.
We will also see that our examples can be interpreted as the result of applying some particular operations to regular and chiral polytopes of high rank. This same operations can be used to obtain alternating semi-regular polyhedra with trivial facet stabilizer (that is, semiregular polyhedra with two face-orbits and where each edge lies between faces in different orbits).
In Section 2 we introduce the basic notions necessary for this paper. Section 3 presents the concept of symmetry type graph, which is used a lot later in the paper. Section 4 presents the non-standard definition of graph used in this paper, together with the concept of voltage graphs and how to construct the derived graph of a voltage graph. In Section 5 we characterize the regular symmetry type graphs that semiregular facet-transitive polyhedra with trivial facet stabilizer may have. Section 6 and Section 7 deal with how to construct polyhedra with the symmetry types determined in Section 5. In Section 8 we interpret our construction in terms of operations. In Section 9 we see how to use these operations to get examples of alternating semiregular polyhedra. Finally, in Section 10 we see how we could generalize some of our constructiions to get semiregular, facet-transitive maniplexes with trivial facet stabilizer in higher rank, although we are not sure yet how to make sure that they are polytopes.
2 Basic concepts
Abstract polytopes, introduced by Schulte and Danzer in [3], are typically defined as posets with certain properties that make them similar to the face lattice of a convex polytope. However, thanks to [4, Theorem 5.3], we know that they can be equivalently defined as a particular kind of maniplex. For simplicity, we will use this other definition.
An $n$-maniplex (or maniplex of rank $n$) is a connected simple graph $\mathcal{M}$ with a coloring of its edges with colors in $[0,n-1]:=\{0,1,\ldots,n-1\}$ such that
  • every vertex is incident to exactly one edge of each color, and
  • given two colors $i$ and $j$ such that $|i-j|>1$, the alternating paths of length 4 using edges of these two colors are closed.
The vertices of a maniplex are referred to as flags. For the purposes of this paper all maniplexes are finite, that is, they have a finite number of flags. Maniplexes are a natural generalization of maps to higher dimmensions. To see other definitions of maniplex and also learn the topological interpretation of this definition see [20].
Given an $n$-maniplex $\mathcal{M}$ and a color $i\in[0,n-1]$, we denote by $\mathcal{M}_{\overline{i}}$ the subgraph of $\mathcal{M}$ obtained by removing the edges of color $i$. The connected components of $\mathcal{M}_{\overline{i}}$ are the $i$-faces of $\mathcal{M}$. The $(n-1)$-faces of an $n$-maniplex are its facets, while the 0-faces and 1-faces are its vertices and edges respectively.
We say that an $i$-face $F$ and a $j$-face $G$ of a maniplex are incident if they have non-empty intersection. If in addition $i\leq j$, we write $F\leq G$. Note that in this way the faces of a maniplex form a poset.
In Figure 1 we see how a hexagon can be interpreted as a 2-maniplex. Note that the vertices of the hexagon correspond to the edges of color 1 of the graph, and the edges of the hexagon correspond to the edges of color 0. Analogously, in Figure 2 we see a triangular prism interpreted as a 3-maniplex. Here the vertices of the prism correspond to cylces with edges of colors 1 and 2 in the graph, the edges of the prism correspond to cycles with edges of colors 0 and 2 in the graph, and the 2-faces of the prism correspond to cycles with edges of colors 0 and 1 in the graph.
Figure 1: A hexagon viewed as a 2-maniplex.
Figure 2: A prism viewed as a 3-maniplex.
Given a maniplex $\mathcal{M}$ we define its dual $\mathcal{M}^{*}$ as the maniplex we get by replacing the edges of color $i$ with edges of color $n-1-i$ for each $i\in[0,n-1]$. Note that the vertices of $\mathcal{M}^{*}$ correspond to the facets of $\mathcal{M}$ and vice-versa. If $\mathcal{M}^{*}$ is isomorphic to $\mathcal{M}$ we say that $\mathcal{M}$ is self-dual.
An abstract $n$-polytope, or abstract polytope of rank $n$, is an $n$-maniplex $\mathcal{P}$ satisfying the following condition:
Path intersection property: Let $\Phi$ and $\Psi$ be two flags of $\mathcal{P}$. If there is a path from $\Phi$ to $\Psi$ that uses only colors in $[0,m]$ and another path from $\Phi$ to $\Psi$ that uses only colors in $[k,n-1]$, then there is a path from $\Phi$ to $\Psi$ that uses only colors in $[k,m]$.
In [4, Theorem 5.3] it is proved that this definition is equivalent to the poset definition given by Shculte and Danzer in  [3]. From here on, whenever we talk about polytopes in this paper we will be thinking about abstract polytopes.
An important property of polytopes when thought of as maniplex is that the flag $\Phi^{i}$, $i$-adjacent to the flag $\Phi$ has the same $j$-face as $\Phi$ for every $j\neq i$, but it has a different $i$-face. In fact, in the study of abstract polytopes this is usually the definition of $i$-adjacent flags.
An isomorphism of $n$-maniplexes is just a graph isomorphism that preserves the colors of the edges. Naturally, an automorphism of a maniplex is just an isomorphism onto itself. The automorphism group of a maniplex $\mathcal{M}$ is denoted by $\Gamma(\mathcal{M})$. It is known that the automorphism group of a maniplex acts freely on its flags (the proof is identical to that of [10, Proposition 2A4], which is the particular case of abstract polytopes). We say that a maniplex $\mathcal{M}$ is regular (some authors prefer reflexive) if $\Gamma(\mathcal{M})$ acts transitively (and therefore regularly) on the flags of $\mathcal{M}$.
A maniplex is called semiregular if it is vertex-transitive (that is, transitive on its 0-faces) and each of its facets is a regular $(n-1)$-maniplex. Note that, since all polygons (2-polytopes) are regular, for polyhedra (3-polytopes) semiregularity is the same as vertex-transitivity.
Given a maniplex $\mathcal{M}$, we define $r_{i}$ as the flag permutation that maps each flag to its $i$-adjacent flag. It is easy to see that $r_{i}$ is an involution and that, if $|i-j|>1$, then $r_{i}r_{j}$ is also an involution. In particular $r_{i}$ and $r_{j}$ commute whenever $i$ and $j$ are not consecutive. If there is an $i\in[0,n-2]$ such that $(r_{i}r_{i+1})^{2}$ fixes some flag, we say that $\mathcal{M}$ is degenerate, otherwise it is non-degenerate.
The group ${\rm Mon}(\mathcal{M})=\langle r_{i}:0\leq i\leq n-1\rangle$ is called the monodromy group of $\mathcal{M}$, and we will call each of its elements a monodromy. It is worth mentioning that some authors prefer the term connection group and the notation $\mathrm{Con}(\mathcal{M})$. An alternative way of defining the automorphism group of a maniplex $\mathcal{M}$ is as the set of flag permutations that commute with the elements of the monodromy group.
For this paper, the automorphism group will act always on the right-hand side and the monodromy group on the left-hand side.
If $\mathcal{M}$ is a regular maniplex, for a fixed flag $\Phi$ we can define $\rho_{i}$ as the (unique) automorphism that maps $\Phi$ to $\Phi^{i}$. It is known that $\{\rho_{i}:i\in\{0,1,\ldots,n-1\}\}$ is a generating set for the automorphism group of $\mathcal{M}$ (see [10, Theorem 2B8] for the polytopal case). Moreover, the function mapping $\rho_{i}$ to $r_{i}$ can be extended to a group isomorphism between $\Gamma(\mathcal{M})$ and ${\rm Mon}(\mathcal{M})$ (see, for example, [13, Theorem 3.9] or [20, Section 7]). In particular, the monodromy group of a regular maniplex acts regularly on its flags. It is also worth noticing that a regular maniplex is degenerate if and only if $\rho_{i}\rho_{i+1}$ is an involution for some $i\in[0,n-2]$.
3 Symmetry type graphs
The symmetry type graph of a polytope (or maniplex) $\mathcal{P}$, denoted by $\mathcal{T}(\mathcal{P})$, is a graph whose vertices are the flag-orbits of $\mathcal{P}$ that has one edge of color $i$ between the orbit of $\Phi$ and the orbit of $\Psi$ if and only if $\Psi$ is in the same orbit as $\Phi^{i}$. If $\Phi$ and $\Phi^{i}$ are in the same orbit, we draw a semi-edge of color $i$ at the vertex representing that orbit in the symmetry type graph.
The symmetry type graph of an $n$-polytope or $n$-maniplex is an example of an $n$-premaniplex, defined as a graph with a coloring of its edges with colors in $\{0,1,\ldots,n-1\}$ such that each vertex is incident to exactly one edge of each color, and satisfying that whenever $|i-j|>1$, the alternating paths of length 4 with colors $i$ and $j$ are closed. For the purposes of this paper all premaniplexes are considered to be connected. In contrast to a maniplex, a premaniplex needs not to be simple. The notions of isomorphism, automorphism, regularity, monodromy, and so on, extend naturallly from maniplexes to premaniplexes.
Given two $n$-premaniplexes $\mathcal{M}$ and $\mathcal{X}$, a covering projection (covering for short) from $\mathcal{M}$ to $\mathcal{X}$, is a function $p$ from the vertices of $\mathcal{M}$ to the vertices of $\mathcal{X}$ that preserves $i$-adjacencies. It can be proved that all coverings are surjective. If $e$ is an edge of $\mathcal{M}$ connecting vertices $x$ and $y$, we define $p(e)$ as the edge connecting $p(x)$ and $p(y)$ that has the same color as $e$. If there is a covering from $\mathcal{M}$ to $\mathcal{X}$, we say that $\mathcal{M}$ covers $\mathcal{X}$. Note that if $\mathcal{M}$ is a maniplex and $\mathcal{X}$ is its symmetry type graph, the natural projection from $\mathcal{M}$ to $\mathcal{X}$ is always a covering.
Some notable symmetry type graphs are the following:
  • The symmetry type graph of regular $n$-polytopes is denoted by ${\mathbf{1}}^{n}$. It has exactly one vertex and semi-edges of each color from 0 to $n-1$ at that vertex.
  • The symmetry type graph of chiral $n$-polytopes is denoted by ${\mathbf{2}}^{n}_{\emptyset}$: It has exactly two vertices and they are joined by one edge of each color from 0 to $n-1$.
  • More in general, if $\mathcal{P}$ is a polytope with two flag orbits and $I\subset\{0,1,\ldots,n-1\}$ is the set of colors such that $\Phi$ and $\Phi^{i}$ are in the same orbit, then the symmetry type graph of $\mathcal{P}$ is ${\mathbf{2}}^{n}_{I}$, which has exactly two vertices joined by one edge of each color in $\overline{I}$ and with semi-edges of each color in $I$ at each vertex.
The symmetry type graph of a polytope can be used to know how many orbits the polytope has in $i$-faces. In fact, if $\mathcal{X}=\mathcal{T}(\mathcal{P})$ then each connected component of $\mathcal{X}_{\overline{i}}$ (the graph obtained from $\mathcal{X}$ by removing the edges of color $i$), corresponds to an orbit of $i$-faces (see [2, Proposition 3.2]). In particular, $\mathcal{P}$ is $i$-face-transitive if and only if $\mathcal{X}_{\overline{i}}$ (the subgraph of $\mathcal{X}$ induced by edges of color different to $i$) is connected.
A polytope is hereditary if every symmetry of each facet induces a symmetry of the whole polytope, or in other words, if the stabilizer of each facet coincedes with its automorphism group. In terms of the symmetry type graph, a polytope $\mathcal{P}$ with symmetry type graph $\mathcal{X}$ is hereditary if each connected component of $\mathcal{X}_{\overline{n-1}}$ is isomorphic to the symmetry type graph of a facet in the corresponding orbit. In particular, if $\mathcal{P}$ has regular facets and is hereditarty, then each connected component of $\mathcal{X}_{\overline{n-1}}$ must have only one vertex, which in turn means that $\mathcal{X}$ is either ${\mathbf{1}}^{n}$ (implying that $\mathcal{P}$ is regular), or ${\mathbf{2}}^{n}_{\overline{n-1}}$: the graph with two vertices joined by an edge of color $n-1$ and semi-edges of all other colors at each vertex (for more on this, see [11]). Note that if we remove edges of any color other than $n-1$ from these graphs we do not disconnect them, meaning that hereditary polytopes are $i$-face transitive for all $i<n-1$. Polytopes with symmetry type graph ${\mathbf{2}}^{n}_{\overline{n-1}}$ are a particular case of alternating semiregular polytopes: semiregular polytopes with two facet orbits where $(n-1)$-adjacent flags are in different orbits. This kind of polytopes have been studied in [14], for example.
In this paper we want to focus on semiregular polytopes that are the opposite of hereditary: Instead of wanting every symmetry of a facet to induce a symmetry of the whole polytope, we want that none of them do, that is, we want the stabilizer of every facet to be trivial. This means that any two flags in the same facet must be in different orbits. In terms of the symmetry type graph, this means that if $\mathcal{P}$ is a polytope with the desired property and $\mathcal{X}$ is its symmetry type graph, each connected component of $\mathcal{X}_{\overline{n-1}}$ must be isomorphic to a facet in the corresponding orbit.
Note that if a hereditary polytope with regular facets is not regular, it must have exactly two facet orbits. This is not the case when we remove hereditarity: semiregular polytopes can, in principle, have any number of facet orbits.
It is also possible to have a facet-transitive semiregular polytope that is not only not regular, but that has as many flag-orbits as the number of flags in a facet. We will focus most of this paper on constructing examples with this exact characteristic.
4 Voltage graphs
Voltage graphs are a tool used in graph theory to construct graphs with a given quotient (see [5]). They have recently been used to construct maniplexes and polytopes with some prescribed symmetry types (see [6] and [12]), as we will do in this paper.
To define voltage graphs first we need to think of graphs in a way that is not completely standard. For this paper, a graph $\mathcal{X}$ consists of a set of vertices $V(\mathcal{X})$, a set of darts $D(\mathcal{X})$, a function $I:D(\mathcal{X})\to V(\mathcal{X})$ and an involution $(\cdot)^{-1}:D(\mathcal{X})\to D(\mathcal{X})$. We call $I(d)$ the initial vertex of $d$ and $d^{-1}$ the inverse of $d$. An edge of $\mathcal{X}$ is just a set of the form $\{d,d^{-1}\}$. A semi-edge is a dart $d$ satisfying $d=d^{-1}$. If $d$ is a semi-edge, we abuse language and call the edge $\{d\}$ a semi-edge as well. We think of the edge $e=\{d,d^{-1}\}$ as a line joining the vertices $I(d)$ and $T(d):=I(d^{-1})$ (called the final vertex of $d$), and we think of its darts as the two (or one) possible orientations. A path $W$ on $\mathcal{X}$ is a finite sequence $d_{1}d_{2}\cdots d_{\ell}$ of darts in $\mathcal{X}$ such that $T(d_{i})=I(d_{i+1})$ for all $i\in\{0,1,\ldots,\ell-1\}$. The vertex $I(d_{1})$ is the initial vertex of $W$ while $T(d_{\ell})$ is its final vertex. The number $\ell$ is called the length of $W$. We refer the reader to [9] for details.
A voltage graph is a triplet $(\mathcal{X},\Gamma,\xi)$ where $\mathcal{X}$ is a graph, $\Gamma$ is a group (called the voltage group), and $\xi:D(\mathcal{X})\to\Gamma$ is a function (called voltage assignment), satisfying that $\xi(d^{-1})=\xi(d)^{-1}$ for every dart $d\in D(\mathcal{X})$. The element $\xi(d)$ is called the voltage of $d$. Moreover, if $W=d_{1}d_{2}d_{3}\cdots d_{k}$ is a path in $\mathcal{X}$, then we define $\xi(W)$ as $\xi(d_{k})\cdots\xi(d_{2})\xi(d_{1})$, and call this the voltage of $W$. Note that if $W$ ends at the initial vertex of $W^{\prime}$, then $\xi(WW^{\prime})=\xi(W^{\prime})\xi(W)$.
Given a voltage graph $(\mathcal{X},\Gamma,\xi)$ we can construct a graph $\mathcal{X}^{\xi}$ where $\Gamma$ acts by automorphisms. This action is regular on both the vertices and darts of $\mathcal{X}$ and the quotient $\mathcal{X}^{\xi}/\Gamma$ is isomorphic to $\mathcal{X}$. To define $\mathcal{X}^{\xi}$ we make $V(\mathcal{X}^{\xi})=V(\mathcal{X})\times\Gamma$ and $D(\mathcal{X}^{\xi})=D(\mathcal{X})\times\Gamma$. Then we define the initial vertex of the dart $(d,\gamma)$ to be $I(d,\gamma)=(I(d),\gamma)$; and the inverse of $(d,\gamma)$ to be $(d,\gamma)^{-1}=(d^{-1},\xi(d)\gamma)$. In other words, if there is a dart $d$ from $x$ to $y$ in $\mathcal{X}$, then there is a dart $(d,\gamma)$ from $(x,\gamma)$ to $(y,\xi(d)\gamma)$ in $\mathcal{X}^{\xi}$. If $\mathcal{X}$ has a coloring (of its vertices, darts and/or edges) we color $(x,\gamma)$ with the same color as $x$.
One can verify that the $\Gamma$-action $(x,\gamma)\tau=(x,\gamma\tau)$ (where $x$ might be either a vertex or a dart) on $\mathcal{X}^{\xi}$ is by (color-preserving) automorphisms, regular and that two elements (darts or vertices) are in the same $\Gamma$-orbit if and only if they have the same first coordinate. Therefore, $\mathcal{X}^{\xi}/\Gamma$ is isomorphic to $\mathcal{X}$. This gives a natural embedding of $\Gamma$ into the automorphism group of $\mathcal{X}^{\xi}$.
It is convinient to work with voltage graphs where $\mathcal{X}$ has a spanning tree $T$ satisfying that all its darts have trivial voltage. In this case, each connected component of $\mathcal{X}^{\xi}$ has $k|V(\mathcal{X})|$ vertices, where $k=|\langle\xi(d):d\in D(\mathcal{X})\setminus D(T)\rangle|$. To be more precise, the connected component of the vertex $(x,\gamma)$ has vertex set $V(\mathcal{X})\times\gamma\langle\xi(d):d\in D(\mathcal{X})\setminus D(T)\rangle$.
A voltage premaniplex is a voltage graph $(\mathcal{X},\Gamma,\xi)$ where $\mathcal{X}$ is a premaniplex and $\xi$ satisfies the following properties:
  • $\Gamma=\langle\xi(d):d\in D(\mathcal{X})\setminus D(T)\rangle$ for a spanning tree $T$ of $\mathcal{X}$ with trivial voltage on all its darts.
  • The voltage of every semiedge has order 2.
  • If $d$ and $d^{\prime}$ are parallel darts (that is, they share their initial and final vertices) then $\xi(d)\neq\xi(d^{\prime})$.
  • Whenever $|i-j|>1$, the voltage of any alternating path of length 4 with colors $i$ and $j$ is trivial.
In [7, Lemma 3.1] it is proven that if $(\mathcal{X},\Gamma,\xi)$ is a voltage premaniplex, then $\mathcal{X}^{\xi}$ is a maniplex. As mentioned before, $\mathcal{X}^{\xi}/\Gamma$ is isomorphic to $\mathcal{X}$, so if one can prove that every automorphism of $\mathcal{X}^{\xi}$ can be interpreted as an element of $\Gamma$, then $\mathcal{X}$ is the symmetry type graph of $\mathcal{X}^{\xi}$.
In general it can be hard to prove that $\Gamma$ is the whole automorphism group of $\mathcal{X}^{\xi}$. However, some automorphisms of $\mathcal{X}^{\xi}$ not in $\Gamma$ are easier to understand than others. An automorphism $\widetilde{\tau}$ of $\mathcal{X}^{\xi}$ is a lift of an automorphism $\tau$ of $\mathcal{X}$ if it acts as $\tau$ on the first coordinate. In this case we also say that $\widetilde{\tau}$ projects (to $\tau$) and that $\tau$ lifts (to $\widetilde{\tau}$). Note that an automorphism of $\mathcal{X}$ may have many lifts (in fact, it has either 0 or $|\Gamma|$ lifts).
The following theorem describes when an automorphism of a voltage graph lifts.
Theorem 4.1.
[9, Theorem 4.1] Let $(X,\Gamma,\xi)$ be a voltage graph such that $\mathcal{X}^{\xi}$ is connected. Let $x$ be a vertex of $\mathcal{X}$ and $\tau$ an automorphism of $X$. Then $\tau$ lifts if and only if the function that maps $\xi(W)$ to $\xi(W\tau)$ for every closed path $W$ based at $x$ is well defined. In this case, this function is an automorphism of $\Gamma$.
A way of interpreting the previous theorem is by saying that an automorphism of $\mathcal{X}$ lifts if and only if it induces an automorphism of $\Gamma$. This is very useful, as we can understand automorphisms that lift in terms of automorphisms of the voltage group. In particular, if we can prove that no non-trivial automorphism of $\mathcal{X}$ induces an automorphism of $\Gamma$, we will know that the automorphisms of $\mathcal{X}^{\xi}$ that project are precisely the elements of $\Gamma$.
The automorphisms that do not project are, in general, much harder to identify. Fortunately, we can avoid them completely in the case when $\mathcal{X}$ is a regular premaniplex, by using the following theorem:
Theorem 4.2.
[1, Corollary 5.2] Let $(\mathcal{X},\Gamma,\xi)$ be a voltage premaniplex. If $\mathcal{X}$ is regular, then every automorphism of $\mathcal{X}^{\xi}$ projects.
5 Regular symmetry type graphs of semiregular facet-transitive polyhedra with trivial facet stabilizer
In this section we will characterize those regular premaniplexes that are symmetry type graps of semiregular, facet-trasnitive polyhedra with trivial facet stabilizer. Recall that, since all abstract polygons are regular, an abstract polyhedron is semiregular if and only if it is vertex-transitive, so we are looking for polyhedra that are both vertex- and facet-transitive.
Let us ask: What could be the symmetry type graph $\mathcal{X}$ of a facet-transitive semiregular ($n+1)$-polytope $\mathcal{P}$ with trivial facet stabilizer? As established in Section 3, $\mathcal{X}_{\overline{n}}$ must be connected for $\mathcal{P}$ to be facet-transitive, and it must be isomorphic to a facet of $\mathcal{P}$ to ensure that facet-stabilizers are trivial in $\mathcal{P}$. Therefore, to construct $\mathcal{X}$ we may start with an $n$-polytope $\mathcal{K}$ (which will be isomorphic to a facet of $\mathcal{P}$), and then add edges of color $n$ in such a way that:
  1. 1.
    The resulting graph $\mathcal{X}$ is an $(n+1)$-premaniplex.
  2. 2.
    $\mathcal{X}_{\overline{0}}$ is connected (to ensure that $\mathcal{P}$ is vertex-transitive).
So, for $n=3$ we are looking for a 3-premaniplex $\mathcal{X}$ satisfying that $\mathcal{X}_{\overline{2}}$ is a polygon, and that $\mathcal{X}_{\overline{0}}$ is connected. In order to be able to use 4.2, we will look for examples where $\mathcal{X}$ is also regular.
Let us find all 3-premaniplexes with those properties:
Start with an $n$-gon, that is, a 2-maniplex with $2n$ flags forming an alternating cycle with colors 0 and 1. We want to add edges of color 2 that ‘‘commute’’ with the edges of color 0 in such a way that our conditions are satisfied. First notice that if there is a semi-edge of color 2, because of regularity, all edges of color 2 would be semi-edges. In this case $\mathcal{X}_{\overline{0}}$ would have $n$ connected components (each consisting of an edge of color 1 and two semi-edges of color 2), and this contradicts our hypothesis that $\mathcal{X}_{\overline{0}}$ is connected. Therefore, there cannot be any semi-edge in $\mathcal{X}$. Since the alternating paths of length 4 with colors 0 and 2 are closed, using regularity we get that either they are all 4-cycles or they are all 2-cycles that are travelled twice. If these paths are 2-cycles, it means that every edge of color 2 is parallel to an edge of color 0. If $\mathcal{X}$ is an $n$-gon with aditional edges of color 2 parallel to its edges of color 0, then $\mathcal{X}_{\overline{0}}$ is an alternating cycle of length $2n$ with colors 1 and 2, which is connected. Therefore, we have found our first family of premaniplexes with the desired properties: $n$-gons with edges of color 2 connecting pairs of 0-adjacent flags (see Figure 3). We will call this family 1.
Figure 3: Symmetry type graphs of family 1.
Now suppose that $\mathcal{X}$ satisfies our conditions and it has a 4-cycle with flags $\Phi$, $\Phi^{0}$, $\Phi^{02}$ and $\Phi^{2}$. There is an automorphism $\tau$ of $\mathcal{X}$ that maps $\Phi$ to $\Phi^{0}$ and $\Phi^{2}$ to $\Phi^{02}$. This means that $\tau$ fixes both the 1-face $e$ that connects $\Phi$ to $\Phi^{0}$ and $e^{\prime}$ that connects $\Phi^{2}$ to $\Phi^{02}$. The only way two different edges of a polygon are fixed by the same automorphism is if they are opposites in a polygon with an even number of sides. So we know that $n$ must be even and that 2-adjacent flags lie on opposite edges. The edge opposite to $e$ has the flags $(r_{1}r_{0})^{n/2}\Phi$ and $r_{0}(r_{1}r_{0})^{n/2}\Phi$, and this edge must be $e^{\prime}$. So, $\Phi^{2}$ must be one of these two flags. If $\Phi^{2}=(r_{1}r_{0})^{n/2}\Phi$ we get an alternating 4-cycle of colors 1 and 2 with flags $\Phi$, $\Phi^{2}$, $\Phi^{21}$ and $\Phi^{1}$. This means that for $n\geq 3$, $\mathcal{X}_{\overline{0}}$ cannot be connected. For $n=2$, however, we have found one sporadic example $\mathcal{X}_{s}$: a digon with edges of color 2 between opposite flags (see Figure 4).
Figure 4: The sporadic premaniplex $\mathcal{X}_{s}$.
Now let us investigate the case when $\Phi^{2}=r_{0}(r_{1}r_{0})^{n/2}\Phi$. Since we are assuming that $\mathcal{X}$ is regular we can use the fact that the monodromy group acts regularly on the flags to get rid of $\Phi$ in our notation and just write $r_{2}=r_{0}(r_{1}r_{0})^{n/2}$. Then, we get that
$(r_{1}r_{2})^{2}=(r_{1}r_{0})^{2(n/2+1)}=(r_{1}r_{0})^{n+2}=(r_{1}r_{0})^{2}.$
Let the edges (1-faces) of $\mathcal{X}$ be called $e_{0},e_{1},\ldots,e_{n-1}$ where $e_{0}$ is the edge of the flag $\Phi$ and the other edges are named cyclically in clockwise order. We will call the flags of the form $(r_{1}r_{0})^{k}\Phi$ left flags, and the flags of the form $r_{0}(r_{1}r_{0})^{k}\Phi$ right flags. Notice that the alternating path of length 4 with flags $\Phi$, $\Phi^{2}$, $\Phi^{21}$ and $\Phi^{212}$ has:
  1. 1.
    the left flag of $e_{0}$,
  2. 2.
    the right flag of $e^{\prime}=e_{n/2}$,
  3. 3.
    the left flag of $e_{n/2+1}$, and
  4. 4.
    the right flag of $e_{1}$.
If we continue the alternating path with colors 1 and 2 we will get
  1. 1.
    the left flags of the edges $e_{2k}$,
  2. 2.
    the right flags of the edges $e_{n/2+2k}$,
  3. 3.
    the left flags of the edges $e_{n/2+2k+1}$, and
  4. 4.
    the right flags of the edges $e_{2k+1}$,
for all $k$.
We want this path to go through every flag before going back to $\Phi$. The only way to make sure that all these flags are different is if $n/2$ is even, so that each 4 steps have a left flag and a right flag of an even edge and of an odd edge. Therefore, our second family of premaniplexes with the desired conditions are $n$-gons where $n$ is divisible by 4, and the edges of color 2 connect each flag to the 0-adjacent flag of its opposite flag (see Figure 5). We will call this family 2.
Figure 5: Symmetry type graphs of family 2.
We have proven the following:
Proposition 5.1.
Let $\mathcal{X}$ be a regular 3-premaniplex such that $\mathcal{X}_{\overline{2}}$ is a polygon and $\mathcal{X}_{\overline{0}}$ is connected. Then $\mathcal{X}$ is isomorphic to either a member of family 1, a member of family 2, or the premaniplex $\mathcal{X}_{s}$.
As discussed before, a semiregular facet-transitive 3-maniplex with trivial facet stabilizer and regular symmetry type graph must have a premaniplex $\mathcal{X}$ satisfying the conditions of 5.1 as its symmetry type graph. However, it is impossible to have a polyhedron $\mathcal{P}$ with trivial facet stabilizer and with $\mathcal{X}_{s}$ as its symmetry type graph: this is because the facets of $\mathcal{P}$ would have to be digons and all the polyhedra with only digonal faces are regular. So we get the following ressult:
Proposition 5.2.
Let $\mathcal{M}$ be a semiregular facet-transitive 3 maniplex with trivial facet stabilizer. Let $\mathcal{X}$ be the symmetry type graph of $\mathcal{M}$. Then $\mathcal{X}$ is isomorphic to a premaniplex in either family 1 or family 2.
6 Constructing polyhedra with symmetry type graph in family 1
Now we want to assign voltages to the premaniplexes of the previous section to get examples of maniplexes and polytopes with these symmetry types. First, let $\mathcal{X}$ be the premaniplex of family 1 with $n$ sides. We want to find a voltage assignment $\xi$ on $\mathcal{X}$ such that $\mathcal{X}^{\xi}$ has symmetry type $\mathcal{X}$ and $n$-gonal faces. To ensure that $\mathcal{X}^{\xi}$ has $n$-gonal faces it is enough to assign trivial voltage to all the edges of color different than 2. Now let $\alpha_{0},\alpha_{1},\ldots,\alpha_{n-1}$ be the voltages of the edges of color 2 in cyclical order. To make sure that $\mathcal{X}^{\xi}$ is a maniplex we only need to make sure that $\alpha_{i}$ has order exactly 2 for every $i$. And finally, to make sure that $\mathcal{X}^{\xi}$ has symmetry type $\mathcal{X}$, we only need to make sure that no automorphism of $\mathcal{X}$ induces an automorphism of our voltage group $\Gamma=\langle\alpha_{0},\ldots,\alpha_{n-1}\rangle$. One way to do this is using the automorphism group of a regular $n$-polytope $\mathcal{P}$ as $\Gamma$ and making $\alpha_{i}=\rho_{i}$, where $\rho_{i}$ is the distinguished generator of $\Gamma(\mathcal{P})$ that maps a base flag $\Phi$ of $\mathcal{P}$ to its $i$-adjacent flag $\Phi^{i}$. If $\mathcal{P}$ is non-degenerate, then the only possible non-trivial permutation of distinguished generators that can induce an automorphism of $\Gamma(\mathcal{P})$ is the one mapping $\rho_{i}$ to $\rho_{n-1-i}$ for every $i$. This permutation corresponds to some reflection of $\mathcal{X}$ and it induces an automorphism of the group if and only if $\mathcal{P}$ is self-dual. So, if $\mathcal{P}$ is not self-dual and non-degenerate, $\mathcal{X}^{\xi}$ is a maniplex with symmetry type $\mathcal{X}$. This voltage assignment has been ilustrated in Figure 6, where the label of an edge of the polygon corresponds to the voltage of the darts of color 2 that start on that edge.
Figure 6: A premaniplex of family 1 with a voltage assignment.
Finally, we would like for $\mathcal{X}^{\xi}$ to be a polyhedron (3-polytope). We will use the following theorem:
Theorem 6.1.
[[1, Proposition 3.14]] Let $(\mathcal{X},\Gamma,\xi)$ be a voltage $(r+1)$-premaniplex. Suppose $\mathcal{X}_{\overline{r}}$ is connected, $\xi(d)=1$ for every dart $d$ of color smaller than $r$ and that $\mathcal{X}^{\xi}$ is a maniplex. Then $\mathcal{X}^{\xi}$ is a polytope if and only if $\mathcal{X}_{\overline{r}}$ is a polytope, and whenever there is a path $W$ with voltage 1 and colors in $[m,r]$, there is also a path $W^{\prime}$ with the same endpoints as $W$ that uses only colors in $[m,r-1]$
In our case $r=2$, so the only interesting case is when $m=1$. In other words, we want to make sure that if two flags are connected by a path with colors 1 and 2 and voltage 1, then they are either the same flag or 1-adjacent flags.
Let us start by analyzing the voltage of an alternating path $W$ of colors 1 and 2 starting from $\Phi$. Suppose $W$ uses a total of $k$ darts of color 2. We can see that if $W$ starts with a dart of color 2, then
$\xi(W)=\prod_{i=0}^{k-1}\rho_{i}$
where the product goes from right to left and the subscripts are taken modulo $n$. On the other hand, if the first dart in $W$ has color 1, then
$\xi(W)=\prod_{i=0}^{k-1}\rho_{n-1-i}.$
However, if $\prod_{i=0}^{n-1}\rho_{i}$ has order $\ell$, we might write this as
$\prod_{i=0}^{k-1}\rho_{n-1-i}=\prod_{i=0}^{n\ell-k-1}\rho_{i}.$
Notice that if the left-hand side has $k$ factors, the right-hand side has $n\ell-k$. Regardless of the color of the first dart, the path $W$ ends in $\Phi$ or $\Phi^{1}$ if and only if $k$ is a multiple of $n$. Therefore, what we want is that $\prod_{i=0}^{k-1}\rho_{i}=1$ if and only if $k$ is a multiple of $n$, and the same is true for $\xi(W)=\prod_{i=0}^{k-1}\rho_{n-1-i}$.
This condition can be interpreted in terms of the Petrie polygon of $\mathcal{P}$. Typically, a Petrie polygon of a polytope is defined as a cyclical sequence of edges (1-faces) such that whenever $12\leq i\leq n-$ 1, any $i$ consecutive edges share an $i$-face but any $i+1$ consecutive edges do not. According to this definition some degenerate polytopes and some non-polytopal maniplexes would not have Petrie polygons, so the following alternative is prefered: We will define a Petrie polygon as a flag orbit under the action of the monodromy $\omega_{\pi}:=r_{n-1}r_{n-2}\cdots r_{0}$.The cardinallity $\ell$ of a Petrie polygon $P$ is called its length. If we look at the sequence of edges $\{(\omega_{\pi}^{i}\Phi)_{1}\}_{i=0}^{\ell}$ on a non-degenerate polytope, we would see that it satisfies the usual definition of Petrie polygon. Actually, in [18] the sequence $\{(\omega_{\pi}^{i}\Phi)_{1}\}_{i=0}^{\ell}$ is called a Petrie scheme and its sequence of edges is called a Petrie polygon, but in this article we do not make that distinction. Notice that the edge $(\omega_{\pi}^{i}\Phi)_{1}$ is incident to the vertices $(\omega_{\pi}^{i}\Phi)_{0}$ and $(\omega_{\pi}^{i+1}\Phi)_{0}$. In this definition we make sure that every flag is in exactly one Petrie polygon. If any two different flags in $P$ have different vertices (0-faces), we will say that $P$ is simple.
The following lemma will be quite useful:
Lemma 6.2.
Let $\mathcal{P}$ be an $n$-polytope. Let $\Phi$ be a flag of $\mathcal{P}$ and let $P$ be the Petrie polygon containing $\Phi$. Let $\ell$ be the length of $P$ and let $S=\{\Psi_{i}\}_{i=0}^{n\ell-1}$ be the sequence of flags defined by $\Psi_{0}=\Phi$ and $\Psi_{i+1}=r_{i}\Psi_{i}$ for $0\leq i\leq n\ell$, where the subscript of $r_{i}$ is taken modulo $n$ and the subscript of $\Psi_{i}$ modulo $n\ell$. If $P$ is simple, then all the elements of $S$ are different.
  • Proof
    Note that $P=\{\Psi_{ni}\}_{i=0}^{\ell-1}$. Notice that whenever $\left\lfloor(j-1)/n\right\rfloor=i-1$ we get that $\Psi_{j}$ has the same vertex (0-face) as $\Psi_{ni}$. Therefore, since $P$ is simple, we get that $\Psi_{i}$ and $\Psi_{j}$ have the same vertex if and only if $\left\lfloor(i-1)/n\right\rfloor=\left\lfloor(j-1)/n\right\rfloor$. As a small caveat, for this statement to be completely true we must rename $\Psi_{0}$ as $\Psi_{n\ell}$.
    Let $i$ and $j$ be such that $\Psi_{i}=\Psi_{j}$ with $10\leq i<j\leq n\ell-$ 1. Then $\Psi_{j}=\prod_{k=i}^{j-1}r_{k}\Psi_{i}$. Since $\Psi_{i}$ and $\Psi_{j}$ have the same vertex, we get that $\left\lfloor(i-1)/n\right\rfloor=\left\lfloor(j-1)/n\right\rfloor$.In particular $j-i<n$, so $r_{i}$ appears exactly once in the product $\prod_{k=i}^{j-1}r_{k}$. But since $\mathcal{P}$ is a polytope, this implies that $\prod_{k=i}^{j-1}r_{k}$ must change the $i-face$ of every flag, so we have reached a contradiction.∎
Going back to the case where $\mathcal{P}$ is a regular $n$-polytope with simple Petrie polygons, using that $\Phi\rho_{i}=r_{i}\Phi$ and the fact that auromorphisms commute with monodromies one can see that
$\Phi\prod_{i=0}^{k}\rho_{i}=\prod_{i=0}^{k}r_{i}\Phi=r_{k}\prod_{i=0}^{k-1}r_{i}\Phi.$
So the sequence $S$ in 6.2 is the same as $\{\Phi\prod_{i=0}^{k-1}\rho_{i}\}_{k=0}^{n\ell-1}$ where the subscripts are taken modulo $n$. Applying the lema we get that $\Phi\prod_{i=0}^{k-1}\rho_{i}=\Phi$ if and only if $k$ is divisible by $n\ell$. In particular, if $\prod_{i=0}^{k-1}\rho_{i}=1$ then $k$ is divisible by $n$.
Now let $W$ be a path with darts of colors 1 and 2. Without loss of generality, we may assume that $W$ is reduced, that is, that it does not have any pair of consecutive inverse darts. This is equivalent to saying that $W$ alternates between colors 1 and 2. Let $W^{\prime}$ be the path that follows the same color sequence as $W$ but starting at the base flag of $\mathcal{X}$. Notice that the voltage of $W$ is a conjugate of the voltage of $W^{\prime}$, so $\xi(W)=1$ if and only if $\xi(W^{\prime})=1$. But we have proved that $\xi(W^{\prime})=1$ if and only if $W^{\prime}$ ends at either the base flag or its 1-adjacent flag, and since $\mathcal{X}$ is regular, the same must be true for $W$.
Therefore, we have arrived to the following result:
Theorem 6.3.
Let $\mathcal{P}$ be a regular $n$-polytope with simple Petrie polygons. Suppose that $\mathcal{P}$ is non-degenerate and non-self-dual. Let $(\mathcal{X},\Gamma,\xi)$ be the voltage premaniplex where $\mathcal{X}$ is an the premaniplex of family 1 with $n$-sides, $\Gamma$ is the automorphism group of $\mathcal{P}$ with distinguished generators $\{\rho_{0},\ldots,\rho_{n-1}\}$, and $\xi$ is the voltage assignment that assigns voltage $\rho_{i}$ to the darts starting at the $i$-th 1-face (as in Figure 6). Then $\mathcal{X}^{\xi}$ is a polyhedron with symmetry type graph isomorphic to $\mathcal{X}$. In particular $\mathcal{X}^{\xi}$ is vertex-transitive, facet-transitive and it has trivial facet stabiliizer.
7 Constructing polytopes with symmetry type graph in family 2
To construct polytopes with symmetry type graphs in family 2 we procede in a similar fashion. Let $\mathcal{X}$ be the premaniplex in family 2 constructed from a $4k$-gon (see Figure 5). We enumerate the 1-faces of $\mathcal{X}$ cyclically starting with $e_{0}$ as the 1-face of the base flag. Note that an edge of color 2 joins a flag in the 1-face $e_{i}$ with one in the edge $e_{2k+i}$ (with the suscripts taken modulo $4k$). We assign trivial voltage to the darts of colors 0 and 1 to ensure that the facets of the derived graph are $4k$-gons. If we assign a voltage $\alpha_{i}$ to a dart going from edge $e_{i}$ to edge $e_{2k+i}$, we must assign the same voltage to the other dart that goes from $e_{i}$ to $e_{2k+i}$. This is to ensure that the derived graph is a maniplex.
To assign the voltages to the darts of color 2, this time we use the automorphism group of a chiral polytope. A maniplex is chiral if it has 2 orbits on flags and adjacent flags are in different orbits. In 2010 Pellicer [15] proved that there are abstract chiral polytopes in every rank greater or equal to 3. Let $\mathcal{P}$ be a chiral polytope of rank $12k+$ 1 and let $\Phi$ be a base flag of $\mathcal{P}$. Because of chirality, for every $i$ in $\{1,2,\ldots,2k\}$, there is an automorphism $\tau_{i}$ that maps $\Phi$ to $r_{i}r_{0}\Phi$. It is worth noting that $\tau_{i}$ is an involution if and only if $i\neq 1$. For $i\in\{0,1,\ldots,2k-1\}$ we will assign the voltage $\tau_{i+1}$ to the darts of color 2 in $\mathcal{X}$ that connect the 1-face $e_{i}$ to $e_{2k+i}$. This voltage assignment is ilustrated in Figure 7, where once again, the labels of the edges of the polygon correspond to the voltages of the darts of color 2 that start on that edge.
Figure 7: A premaniplex of family 2 with a voltage assignment.
Using the fact that $\tau_{1}$ is the only $\tau_{i}$ that is not an involution we can see that the only possible automorphisms of $\mathcal{X}$ that could induce an automorphism of $\Gamma(\mathcal{P})$ are the identiy and the half turn. However, it is known (see [15, Proposition 2.2], for example) that if $\mathcal{P}$ is a chiral polytope, there is no automorphism of $\Gamma(\mathcal{P})$ that maps $\tau_{1}$ to $\tau_{1}^{-1}$ and fixes every other $\tau_{i}$. Therefore, using 4.1 we get that no automorphism of $\mathcal{X}$ lifts, and therefore $\mathcal{X}$ is the symmetry type graph of $\mathcal{X}^{\xi}$.
If we look at a path $W$ of alternating colors 2 and 1 that starts on the left flag of $e_{0}$ and uses $j$ edges of color 2, we will see that its voltage is $\prod_{i=1}^{j}\tau_{i}^{\epsilon_{i}}$, where the subscript of $\tau_{i}$ is taken modulo $2k$ (using $\tau_{2k}$ when $i\equiv 0\mod 2k$) and $\epsilon_{i}$ is 1 unless $i\equiv 2k+1\mod 4k$, in which case it is $-1$. The path $W$ ends in $\Phi$ or $\Phi^{1}$ whenever it uses a multiple of $4k$ edges of color 2. So using 6.1, we get that in order to find examples where the derived graph is polytopal, we need to make sure that $\prod_{i=1}^{j}\tau_{i}^{\epsilon_{i}}=1$ implies that $j$ is divisible by $4k$.
Proposition 7.1.
Let $\mathcal{P}$ be a chiral $(2k+1)$-polytope with a base flag $\Phi$, and let $\tau_{i}$ be such that $\Phi\tau_{i}=r_{i}r_{0}\Phi$ for $1\leq i\leq 2k$ and let $\epsilon_{i}$ be defined as above. Let $j$ be a positive integer. If $j\not\equiv\left\lfloor(j-1)/(2k)\right\rfloor\mod 2$, then
$\Phi\prod_{i=1}^{j}\tau_{i}^{\epsilon_{i}}=\prod_{i=0}^{j+\left\lfloor(j-1)/(2k)\right\rfloor}r_{i}\Phi,$
where the subscript of $r_{i}$ is taken modulo $12k+$ 1 and the subscript of $\tau_{i}$ modulo $2k$. On the other hand, if $j\equiv\left\lfloor(j-1)/(2k)\right\rfloor\mod 2$, then
$\Phi\prod_{i=1}^{j}\tau_{i}^{\epsilon_{i}}=r_{0}\prod_{i=0}^{j+\left\lfloor(j-1)/(2k)\right\rfloor}r_{i}\Phi.$
  • Proof
    We will prove this in a ‘‘hand-wavy’’ manner. The reader may formalize this proof using induction and making several cases. For $j=1$ we get $\Phi\tau_{1}=r_{1}r_{0}\Phi$ by definition of $\tau_{1}$. Now suppose the result is true for $j-1$. Using the definition of $\tau_{j}$ and the fact that automorphisms commute with monodromies we get
    $\Phi\prod_{i=1}^{j}\tau_{i}^{\epsilon_{i}}$ $=$ $\Phi\tau_{j}^{\epsilon_{j}}\prod_{i=1}^{j-1}\tau_{i}^{\epsilon_{i}}$
    $=$ $((r_{j}r_{0})^{\epsilon_{j}}\Phi)\prod_{i=1}^{j-1}\tau_{i}^{\epsilon_{i}}$
    $=$ $(r_{j}r_{0})^{\epsilon_{j}}(\Phi\prod_{i=1}^{j-1}\tau_{i}^{\epsilon_{i}}).$
    Recall that the subscripts of $\tau_{i}$ and $\tau_{j}$ are taken modulo $2k$, and therefore, the subscript $j$ of $r_{j}$ is also taken modulo $2k$. If $j\not\equiv 1\mod 2k$, we see that the exponent $\epsilon_{j}$ is irrelevant and the $r_{0}$ gets canceled or not depending on wether $\Phi\prod_{i=1}^{j-1}\tau_{i}^{\epsilon_{i}}$ has an $r_{0}$ on the left or not. Notice that for $j<2k+1$ there is alternation between these two cases, where the even terms are the ones with an $r_{0}$ on the left. In particular $\Phi\prod_{i=1}^{2k}\tau_{i}^{\epsilon_{i}}=r_{0}\prod_{i=0}^{2k}r_{i}\Phi$. Then for $j=2k+1$ we get
    $\Phi\prod_{i=1}^{2k+1}\tau_{i}^{\epsilon_{i}}$ $=$ $(r_{1}r_{0})^{\epsilon_{2k+1}}(\Phi\prod_{i=1}^{2k}\tau_{i}^{\epsilon_{i}})$
    $=$ $r_{0}r_{1}(r_{0}\prod_{i=0}^{2k}r_{i}\Phi)$
    $=$ $r_{0}\prod_{i=0}^{2k+2}r_{i}\Phi.$
    Notice how at this point the alternation pauses momentarily and we get two consecutive products with an $r_{0}$ on the left. Notice also that the superscript of the product increases by 2 rather than 1. After this point alternation resumes, but now the odd terms are the ones with an $r_{0}$ on the left. In particular $\Phi\prod_{i=1}^{4k}\tau_{i}^{\epsilon_{i}}=\prod_{i=0}^{4k+1}r_{i}\Phi$. Now for $j=4k+1$ we get:
    $\Phi\prod_{i=1}^{4k+1}\tau_{i}^{\epsilon_{i}}$ $=$ $(r_{1}r_{0})^{\epsilon_{4k+1}}(\Phi\prod_{i=1}^{4k+1}\tau_{i}^{\epsilon_{i}})$
    $=$ $r_{1}r_{0}(\prod_{i=0}^{4k+1}r_{i}\Phi)$
    $=$ $\prod_{i=0}^{4k+3}r_{i}\Phi.$
    This time we get two consecutive terms without an $r_{0}$ on the left and once again the superscript has increased by 2. This patterns repeats periodically: usually we alternate between terms with and without $r_{0}$ on the left and we increase the superscript by 1, but when we get to a term congruent with 1 mod $2k$ the alternation stops and the superscript increases by 2.∎
Let $\mathcal{P}$ be a chiral polytope of rank $12k+$ 1 with a base flag $\Phi$ with simple Petrie polygons. Similarily to the case for family 1, we take a path $W$ in $\mathcal{X}$ starting from the base flag, alternating darts of color 2 and 1, starting with color 1 and using a total of $j$ darts of color 2. Using 7.1 we see that the voltage $W$ acts as a $2t$-step rotation in the Petrie polygon $P$ whenever $W$ uses $4kt$ edges of color 2.
Suppose that $\xi(W)=1$, in particular $\Phi=\Phi\xi(W)$. We want to prove that $j$, the number of edges of color 2 used by $W$, is a multiple of $4k$.
If $j\not\equiv\left\lfloor(j-1)/(2k)\right\rfloor\mod 2$, then $\Phi\xi(W)=\prod_{i=0}^{j+\left\lfloor(j-1)/(2k)\right\rfloor}r_{i}\Phi$. Let $t=j+\left\lfloor(j-1)/(2k)\right\rfloor+1$. We can use the same argument as we used for family 1 to prove that $\Phi=\prod_{i=0}^{t-1}r_{i}\Phi$ implies that $t$ is divisible by the rank, which in this case is $12k+$ 1. A simple substitution shows that when $j=2km$ for some positive $m$, then $t=(2k+1)m$. Since $t$ is defined by a monotonous function on $j$, the converse is also true. This shows that if $\prod_{i=0}^{t-1}r_{i}$ fixes the base flag, then $j$ is divisible by $2k$. Since $j\not\equiv\left\lfloor(j-1)/2k\right\rfloor\mod 2$, then $j/(2k)\equiv j\equiv 0\mod 2$, and therefore $j$ must be a multiple of $4k$.
Now suppose $j\equiv\left\lfloor(j-1)/2\right\rfloor\mod 2$ and $\Phi=\Phi\xi(W)=r_{0}\prod_{i=0}^{j+\left\lfloor(j-1)/(2k)\right\rfloor}r_{i}\Phi$. Then $r_{0}\Phi=\prod_{i=0}^{j+\left\lfloor(j-1)/(2k)\right\rfloor}r_{i}\Phi$. Let $t=j+\left\lfloor(j-1)/(2k)\right\rfloor+1$ and let $t^{\prime}$ be its residue modulo $(2k+1)\ell$, where $\ell$ is the length of $P$. Then, by definition of $\ell$ we see that $r_{0}\Phi=\prod_{i=0}^{t-1}\Phi=\prod_{i=0}^{t^{\prime}-1}r_{i}\Phi$. 6.2 tells us that $t^{\prime}=1$, in particular $t\equiv 1\mod 2k+1$. We have established that $t$ is divisible by $12k+$ 1 if and only if $j$ is divisible by $2k$. If $j\equiv r\mod 2k$, for $1\leq r<2k$, a substitution shows that $t\equiv r+1\mod 2k+1$. Therefore, it is not possible that $t\equiv 1\mod 2k+1$. This proves that this case cannot happen.
We have shown that $\xi(W)=1$ implies that $W$ ends at the base flag, or at its 1-adjacent flag. In analogy to the case for family 1, we can use this to prove that any path with colors 1 and 2 with trivial voltage must end either on the flag where it started or on its 1-adjacent flag. This proves the following theorem:
Theorem 7.2.
Let $\mathcal{P}$ be a chiral $n$-polytope with simple Petrie polygons. Suppose that $\mathcal{P}$ is non-degenerate and non-self-dual. Let $(\mathcal{X},\Gamma,\xi)$ be the voltage premaniplex where $\mathcal{X}$ is the premaniplex of family 2 with $4k$ sides, $\Gamma$ is the automorphism group of $\mathcal{P}$ with distinguished generators $\{\tau_{1},\ldots,\tau_{2k}\}$, and $\xi$ is the voltage assignment that assigns voltage $\tau_{i+1}$ to the darts starting at the $i$-th 1-face for $i<2k$ (as in Figure 7). Then $\mathcal{X}^{\xi}$ is a polyhedron with symmetry type graph isomorphic to $\mathcal{X}$. In particular $\mathcal{X}^{\xi}$ is vertex-transitive, facet-transitive and it has trivial facet stabiliizer.
8 Interpreting with voltage operations
In this section we want to interpret the constructions of the previous two sections as applying some well defined operation to a regular or chiral polytope. To do this we will use voltage operations. Voltage operations were introduced in [8], and they are a family of operations that include the Whytoffian constructions, duality, Petrie duality, and even some operations that modify the rank, such as the prism and the pyramid. They are very useful, because they can be applied not only to maniplexes and polytopes, but also to premaniplexes. In fact, if $\mathcal{O}$ is a voltage operation and $\mathcal{M}$ is a maniplex with symmetry type graph $\mathcal{X}$, the symmetry type graph of $\mathcal{O}(\mathcal{M})$ is a quotient of $\mathcal{O}(\mathcal{X})$ (see [8, Theorem 5.1]).
An $(n,m)$-voltage operator is a voltage premaniplex $(\mathcal{Y},W^{n},\eta)$ where $\mathcal{Y}$ has rank $m$ and $W^{n}$ is the group with presentation $\langle r_{0},r_{1},\ldots r_{n-1}\mid r_{i}^{2}=(r_{i}r_{j})^{2}=1\text{ whenever }|i-j|>1\rangle$. Since the monodromy group of any $n$-premaniplex is a quotient of $W^{n}$, this group acts naturally on the set of flags of any $n$-premaniplex (on the left) by monodromies. Note that we are abusing notation by using $r_{i}$ both for an element of $W^{n}$ and an element of ${\rm Mon}(\mathcal{X})$ where $\mathcal{X}$ is a premaniplex, however this is not a problem because both elements act the same way on each flag of $\mathcal{X}$.
Given an $n$-premaniplex $\mathcal{X}$ and an $(n,m)$-voltage operator $(\mathcal{Y},W^{n},\eta)$ we can create a new $m$-premaniplex denoted by $\mathcal{X}\rtimes_{\eta}\mathcal{Y}$ as follows: The flags of $\mathcal{X}\rtimes_{\eta}\mathcal{Y}$ are the cartesian product $\mathcal{X}\times\mathcal{Y}$, and the $i$-adjacent flag of $(x,y)$ is given by
$r_{i}(x,y)=(\eta(^{i}y)x,r_{i}(y)),$
where ${}^{i}y$ is the dart of color $i$ that starts at $y$.
Given a voltage $n$-premaniplex $(\mathcal{X},\Gamma,\xi)$ and a voltage operator $(\mathcal{Y},W^{n},\eta)$,[8, Theorem 5.4] gives us a way to construct a voltage assignment $\theta$ in $\mathcal{X}\rtimes_{\eta}\mathcal{Y}$ with values in $\Gamma$ such that $(\mathcal{X}\rtimes_{\eta}\mathcal{Y})^{\theta}$ is isomorphic to $\mathcal{X}^{\xi}\rtimes_{\eta}\mathcal{Y}$. As a corollary we get
Corollary 8.1.
Let $\mathcal{M}$ be a regular maniplex with automorphism group $\langle\rho_{0},\rho_{1},\ldots,\rho_{n-1}\rangle$ and let $(\mathcal{Y},W^{n},\eta)$ be a voltage operator. Then $\mathcal{X}\rtimes_{\eta}\mathcal{Y}$ is isomorphic to $\mathcal{Y}^{\nu}$ where $\nu$ is the voltage assignment on $\mathcal{Y}$ we obtain from $\eta$ by replacing every occurence of each $r_{i}$ with the corresopnding $\rho_{i}$.
8.1 Family 1
Let $\mathcal{X}$ be the member of family 1 obtained from an $n$-gon. 6.3 tells us that we can construct a polytope whose symmetry type is $\mathcal{X}$ by using the voltage premaniplex in Figure 6, where $\langle\rho_{0},\ldots,\rho_{n-1}\rangle$ is the automorphism group of some non-degenerate, non-self-dual regular $n$-polytope $\mathcal{P}$ with simple Petrie polygons. Then, using 8.1 we get that $\mathcal{X}^{\xi}$ is isomorphic to $\mathcal{P}\rtimes_{\eta}\mathcal{X}$ where $(\mathcal{X},W^{n},\eta)$ is the voltage operation we get by replacing the voltage $\rho_{i}$ by $r_{i}$ (see Figure 8).
Figure 8: The voltage operator $(\mathcal{X},W^{n},\eta)$.
In what follows we want to give some topological interpretation of what the voltage operation $\mathcal{O}:\mathcal{P}\mapsto\mathcal{P}\rtimes_{\eta}\mathcal{X}$ does. First, we need to introduce more concepts. Given a polyhedron $\mathcal{P}$ with simple Petrie-polygons, we define the petrial or Petrie dual of $\mathcal{P}$, as the polyhedron $\mathcal{P}^{\pi}$ with the same vertices and edges, but where the 2-faces correspond to the Petrie polygons of $\mathcal{P}$. We can extend this definition to all 3-maniplexes (and pre-maniplexes) as follows: Given a 3-premaniplex $\mathcal{X}$, its petrial $\mathcal{X}^{\pi}$ is defined as the premaniplex we get by removing the edges of color 0, and adding a new edge of color 0 between each flag $\Phi$ and its $02$-adjacent flag $r_{2}r_{0}\Phi$. If we denote by $\mathcal{X}^{*}$ the usual dual of $\mathcal{X}$, then $((\mathcal{X}^{*})^{\pi})^{*}$ is denoted by $\mathcal{X}^{\rm opp}$ and is known as the opposite of $\mathcal{X}$. It is easy to see that the operations dual, petrial and opposite are voltage operations. The corresponding voltage operators are ilustrated in Figure 9.
Figure 9: The voltage operators for the dual, petrial, and opposite operations.
Using [8, Theorem 6.1], we notice that we can interpret $\mathcal{O}$ as the result of concatenating a different operation $\mathcal{O}^{\prime}$ with the opposite operation. The operation $\mathcal{O}^{\prime}$ is given by the voltage operator $(\mathcal{Y},W^{n},\eta_{1})$, obtained by replacing the edges of color 2 in $\mathcal{X}$ by semi-edges (see Figure 10).
Figure 10: The voltage operator $(\mathcal{Y},W^{n},\eta_{1})$.
If we want to realize an $n$-maniplex as a topological object, what we usually do is interpret each flag $\Phi$ as an $(n-1)$-simplex whose vertices are the baricenters of the faces containing that flag. We can label the vertices of this simplex according to the dimension of the face they are the baricenter of, so that the vertex labeled 0 corresponds to the vertex $\Phi_{0}$, the vertex labeled 1 corresponds to the baricenter of the edge $\Phi_{1}$, etc. Then we can number each facet of the simplex according to its opposite vertex. This means that two simplex sharing a facet labeled $i$ correspond to $i$-adjacent flags.
Figure 10 is telling us that to apply the operation $\mathcal{O}^{\prime}$ we have to replace each flag by an $n$-gon, and that $i$-adjacent flags correspond to $n$-gons sharing their $i$-th side. One way to do this topologically is as follows: We start by representing each flag as a simplex, as usual. Then, for each $i\in\{0,1,\ldots,n-1\}$ we draw a point $u_{i}$ at the baricenter of the intersection of the facets numbered $i$ and $i-1$ (reduced modulo $n$ if necessary). Then we construct an $n$-gon by drawing a sraight segment between $u_{i}$ and $u_{i+1}$ for each $i$. Notice that the edge joining $u_{i}$ to $u_{i+1}$ is contained in the facet with the label $i$. Therefore, the polygons corresponding to a pair of $i$-adjacent flags will share their $i$-th sides, just as requested.
For $n=3$, the intersection of the facet $i$ and the facet $i-1$ is exactly the vertex $i+1$, so the triangle we get by applying $\mathcal{O}^{\prime}$ is the same as the triangle representing the flag. In other words, in rank 3, the operation $\mathcal{O}^{\prime}$ is the baricentric subdivision. In rank 4 the polygons we get have a vertex in the midpoint of each edge of a particular Petrie polygon of the simplex. In Figure 11 we see an example in rank 4. Here the facets of our 4-polytope are cubes, and we illustrate the simplex corresponding to two 0-adjacent flags together with the polygons that they turn into when we apply the operation $\mathcal{O}^{\prime}$.
Figure 11: The operation $\mathcal{O}^{\prime}$ in rank 4.
Note that these operations can be applied to any pre-maniplex of the correct rank, but we can only be certain that the result has symmetry type in family 1 when the operated object is a regular polyhedron that is non-degenerate, non-self dual, and has simple Petrie polygons.
8.2 Family 2
Now let us consider the construction of 7.2. Remeber that this construction starts with a chiral polytope of rank $12k+$ 1 and gives us a polyhedron with $4k$-gonal facets. Using [8, Theorem 5.4] we can prove that the voltage premaniplex in Figure 7 can be obtained by applying a voltage operation $\mathcal{O}$ to the symmetry type graph of the chiral polytope $\mathcal{P}$. The voltage operator corresponding to $\mathcal{O}$ is ilustrated in Figure 12. It is similar to the operation in Figure 8, but the voltages of the edges of color 2 go from $r_{1}$ to $r_{k}$ and the edge of color 1 between the first and last flags in the cycle, has voltage $r_{0}$. We have removed the polygon in gray to keep the image cleaner.
Figure 12: The voltage operator for family 2.
Again, this operation can be interpreted as the concatenation of another operation $\mathcal{O}^{\prime}$ and the opposite operation, which turns the edges of color 2 into semi-edges (see Figure 13).
Figure 13: The operator for $\mathcal{O}^{\prime}$.
For family 2 we now give a topologic interpretation for the operation $\mathcal{O}^{\prime}$ when it turns a 3-polytope (a polyhedron) into a polyhedron with square faces. The corresponding voltage operator appears in Figure 14.
Figure 14: The operator for $\mathcal{O}^{\prime}$ in rank 3.
One can check that the operation $\mathcal{O}^{\prime}$ in rank 3 turns every pair of 0-adjacent flags in $\mathcal{P}$ into a square. Such a pair determines an edge (1-face) and a facet of $\mathcal{P}$. Pairs of 0-adjacent flags that share a 1-face are turned into squares that share two opposite sides, while pairs that have 0-adjacent or 1-adjacent flags turn into squares that share 1 side.
A topological way to do this is as follows: Replace each edge $e$ of $\mathcal{P}$ with a cylinder split in half by a plane containing its axis. This divides the cylinder into two squares (making the ‘‘walls’’ of a digonal prism). We can label each square with one of the two facets of $\mathcal{P}$ containing $e$. One can also label each end of the cylinder with one of the vertices incident to $e$. Then, given a flag $\Phi$ of $\mathcal{P}$ there is a semicircle that corresponds to taking the cylinder corresponding to $\Phi_{1}$, taking the half labeled by $\Phi_{2}$, and taking the side of that square that corresponds to $\Phi_{0}$. Then for each flag $\Phi$ we glue the semicircle corresponding to $\Phi$ to the one corresponding to $r_{1}\Phi$. The resulting polyhedron is $\mathcal{O}^{\prime}(\mathcal{P})$.
We can interpret this operation also in the context of incidence geometries. The vertices of $\mathcal{O}^{\prime}(\mathcal{P})$ can be thought of as $\mathcal{P}_{0}\times\{0,1\}$ where $\mathcal{P}_{0}$ is the vertex-set of $\mathcal{P}$. The edges of $\mathcal{O}^{\prime}(\mathcal{P})$ can be thought of as $\mathcal{P}_{1}\times\{0,1\}$ where $\mathcal{P}_{1}$ is the edge-set of $\mathcal{P}$, together with the pairs of the form $(v,f)$ where $v$ is a vertex of $\mathcal{P}$ and $f$ is a facet of $\mathcal{P}$ incident to $v$. The edge $(e,\varepsilon)$ is incident to the vertex $(v,\delta)$ if and only if $\varepsilon=\delta$ and $e$ is incident to $v$ in $\mathcal{P}$, while the edge $(v,f)$ is incident to $(v,0)$ and $(v,1)$. The facets of $\mathcal{O}^{\prime}(\mathcal{P})$ correspond to pairs $(e,f)$ where $e$ is an edge of $\mathcal{P}$ and $f$ is a facet of $\mathcal{P}$ incident to $e$. The facet $(e,f)$ is incident to the edges $(e,0)$, $(e,1)$, $(v,f)$ and $(u,f)$, where $v$ and $u$ are the vertices of $\mathcal{P}$ incident to $e$ (see Figure 15).
Figure 15: The operation $\mathcal{O}^{\prime}$ in rank 3.
One can verify that the voltage operator corresponding to this operation is exactly the one in Figure 14.
Once again, the operations described in this section can be applied to all premaniplexes provided they have the correct rank, but we are only guaranteed to get a polytope with symmetry type in family 2 when the operated object is a chiral polytope with simple Petrie polygons.
9 Alternating semiregular polyhedra with trivial facet stabilizer
In Section 8.1 we have discussed how polyhedra with symmetry type graph in family 1 can be obtained by applying the voltage operation illustrated in Figure 8 to a regular polytope of rank $n$. If instead we apply this same operation to a chiral polytope $\mathcal{P}$ with automorphism group $\Gamma(\mathcal{P})=\langle\tau_{1},\tau_{2},\ldots,\tau_{n-1}\rangle$, where once again $\Phi\tau_{i}=r_{i}r_{0}\Phi$ for all $i$ and some fixed base flag $\Phi$, then [8, Theorem 6.1] tells us that the result is the derived graph of the voltage graph in Figure 16. This graph is obtained by taking two $n$-gons and then connecting corresponding edges with a ‘‘cross’’ of edges of color 2. That is, the left flag of the $i$-th edge of one copy is 2-adjacent to the right flag of the $i$-th edge of the second copy and viceversa. The darts of color $2$ that start on the $i$-th edge first copy have voltage $\tau_{i}$, and the voltage of the darts of color 2 starting at the 0-th edge of any copy have trivial voltage. The way to construct this graph is to apply the voltage operation in question to the symmetry type graph of a chiral $n$-polytope, which is $2^{n}_{\emptyset}$. See [8, Theorem 6.1] to see how to determine the voltage assignment.
Figure 16: The voltage graph resulting from applying the operation of Section 8.1 to $2^{n}_{\emptyset}$.
Let $(\mathcal{X},\Gamma(\mathcal{P}),\xi)$ be the voltage graph in Figure 16. Observing that only the edges of color two can have non-trivial voltage we can see that the facets of the derived graph will all be $n$-gons. If we remove the edges of color 2 from $\mathcal{X}$ we get two connected components. This means that $\Gamma(\mathcal{P})$ acts with two orbits on the set of facets. On the other hand, the number of connected components after removing the edges of color 0 depends on the parity of $n$: for odd $n$ we get one connected component, while for even $n$ we get two connected components.
Again, by observing that no non-trivial permutation of the set $\{\tau_{i}^{\pm 1}:1\leq i\leq n-1\}$ induces an automorphism of $\Gamma(\mathcal{P})$ we can see that no automorphism of $\mathcal{X}$ lifts, and therefore $\mathcal{X}$ is the symmetry type graph of $\mathcal{X}^{\xi}$.
Notice also that, since every edge of color 2 connects the two components of $\mathcal{X}_{\overline{2}}$, we know that every 1-face (edge) of $\mathcal{X}^{\xi}$ is shared by facets in different orbits.
Therefore, if $n$ is odd then $\mathcal{X}^{\xi}$ is an alternating semiregular maniplex with trivial facet stabilizer. Now we want to prove that it is also polytopal.
To do this we will use the following generalization of 6.1 (the proof is identical):
Theorem 9.1.
Let $(\mathcal{X},\Gamma,\xi)$ be a voltage $(r+1)$-premaniplex. Suppose $\xi(d)=1$ for every dart $d$ of color smaller than $r$. Then $\mathcal{X}^{\xi}$ is a polytope if and only if the connected components of $\mathcal{X}_{\overline{r}}$ are polytopes, and whenever there is a path $W$ with voltage 1 and colors in $[m,r]$ ending in the same connected component of $\mathcal{X}_{\overline{r}}$ as it started, there is also a path $W^{\prime}$ with the same endpoints as $W$ that uses only colors in $[m,r-1]$.
Let us denote by $\phi$ (lowercase) the right flag of the outer $n$-gon. If we start a path $W$ at $\phi$ and follow a path that alternates between colors 1 and 2 (starting with color 1), we will see that its voltage is $\prod_{i=1}^{j}\tau_{i}^{\epsilon_{i}}$, where the subscript of $\tau_{i}$ is taken modulo $n-1$ (using $\tau_{n-1}$ when $i\equiv 0\mod n-1$), $\epsilon_{i}$ is $-1$ whenever $i\equiv n+1\mod 2n$ and 1 in any other case and $j$ is the number of edges of color 2 with non-trivial voltage used by $W$. By using 7.1 we get that the voltage of $W$ acts like a $2t$-step rotation in the Petrie poligon of $\Phi$ when $j=2(n-1)t$. Using the same argument as in 7.2 we get that if $W$ has trivial voltage, then $j$ is divisible by $2(n-1)$. This means that the last dart with non-trivial voltage used by $W$ is the dart of color 2 ending on the right flag of the $(n-1)$-th edge of the inner $n$-gon. (which is $r_{1}r_{2}\phi$). Therefore, $W$ must end in either $r_{1}r_{2}\phi$, $r_{2}\phi$, $\phi$ or $r_{1}\phi$, but only the last two are in the same connected component of $\mathcal{X}_{\overline{2}}$ as $\phi$. So, by using 9.1 we get the following result.
Theorem 9.2.
Let $\mathcal{P}$ be a chiral $n$-polytope with simple Petrie polygons. Let $\mathcal{O}$ be the voltage operation of Section 8.1. Then $\mathcal{O}(\mathcal{P})$ is an alternating semiregular polyhedron with trivial facet stabilizer and its symmetry type graph is the one in Figure 16.
10 Higher rank
We can generalize family 1 to construct some examples of semiregular maniplexes with trivial facet stabilizer and exactly one orbit on facets.
Let $\mathcal{M}$ be a regular $n$-maniplex with facets isomorphic to an $(n-1)$-maniplex $\mathcal{K}$. Suppose that $\Gamma(\mathcal{K})$ has an involution $\sigma$ in its center. Given a flag $\Phi$ in a facet $F$ of $\mathcal{M}$ we define $r_{n}\Phi$ as $\Phi\tau_{F}^{-1}\sigma\tau_{F}$, where $\tau_{F}$ is some isomorphism from $\mathcal{K}$ to $F$. One can prove that $r_{n}$ does not depend on the choice of $\tau_{F}$ (however, it does depend on the choice of $\sigma$). We may call $r_{n}\Phi$ the opposite flag of $\Phi$ in the facet $F$. Moreover, if $G$ is an $i$-face incident to $F$ with $i<n-1$, we can define the opposite face of $G$ in $F$ in a natural way. By adding edges of color $n$ between each flag and its opposite (in its respective facet), we get an $(n+1)$-premaniplex $\mathcal{X}$ (in fact, it is usually a maniplex). If $(\mathcal{X})_{\overline{0}}$ is connected, then we can use $\mathcal{X}$ to construct a semiregular maniplex with symmetry type $\mathcal{X}$:
Let $F_{0},F_{1},\ldots,F_{N-1}$ be the facets of $\mathcal{M}$. Let $\mathcal{P}$ be a regular polytope of rank $N$. Let $\Gamma(\mathcal{P})=\langle\rho_{0},\ldots,\rho_{N-1}\rangle$. We construct a voltage assignment $\xi:D(\mathcal{X})\to\Gamma(\mathcal{P})$ by giving trivial voltage to all darts of color smaller than $n$, and giving voltage $\rho_{i}$ to the darts of color $n$ in the facet $F_{i}$. Then $\mathcal{X}^{\xi}$ will be an $(n+1)$-maniplex whose facets are isomorphic to $\mathcal{M}$. If no automorphism of $\mathcal{X}$ induces an automorphism of $\Gamma(\mathcal{P})$, we can be sure that $\mathcal{X}$ is the symmetry type graph of $\mathcal{X}^{\xi}$, immplying that $\mathcal{X}^{\xi}$ has trivial facet stabilizer. If $\mathcal{P}$ is non-degenerate, the only possible non-trivial automorphism induced by a permutation of the generators of $\mathcal{P}$ would be one that maps $\rho_{i}$ to $\rho_{N-1-i}$. In most cases, we will be able to find a labeling of the facets of $\mathcal{M}$ in a way that there is no automorphism that induces this permutation. In the rare case that this happens to be impossible, by demanding $\mathcal{P}$ to not be self-dual we make sure that this permutation of generators does not induce an automorphism of $\Gamma(\mathcal{P})$.
Just as we did in Section 8.1, we may interpret this construction as applying some voltage operation that turns regular $N$-polytopes into semiregular, facet-transitive $(n+1)$-maniplexes with trivial facet stabilizer. Applying this operation to a chiral $N$-polytope would give an alternating semiregular $(n+1)$-maniplex with trivial facet stabilizer (just as in Section 9.
The problem comes with polytopality. We can be certain that we need $\mathcal{M}$ to be a polytope for $\mathcal{X}^{\xi}$ to be a polytope. However it is not obvious what other conditions are necessary or sufficient to make sure that $\mathcal{X}^{\xi}$ is a polytope. This is an interesting problem to explore in the future. We have tried doing some experimentation with RAMP [19], but the examples become too big too quickly.
Acknowledgments
This paper was completed while the author had a postdoctoral position at the Department of Mathematics of Northeastern University.